A chemist has a beaker of a 3% acid solution and a beaker of a 7% acid solution.He needs to make 75mL of a 4% acid solution.With the information given,write a system of linear equations and solve the system.
Answer by stanbon(75887)

A chemist has a beaker of a 3% acid solution and a beaker of a 7% acid solution.He needs to make 75mL of a 4% acid solution.
-----
Quantity: t + s = 75
Acid:::: 3t +7s = 4*75
----
Multiply thru the Quantity equation by 3 to get:
3t + 3s = 75
3t + 7s = 4*75
---
Subtract and solve for "s":
4s = 3*75
s = 56.25 oz (amt. of 7% solution needed)
--------------------------
Solve for "t":
t + 56.25 = 75
t = 18.75 oz (amt. of 3% solution needed)
==============================================
Cheers,
Stan H.
==============